# Neural Topic Modeling with large language models in the loop

### v1

@@ Abstract
Topic modeling is a fundamental task in natural language processing, allowing the discovery of latent thematic structures in text corpora. While Large Language Models (LLMs) have demonstrated promising capabilities in topic discovery, their direct application to topic modeling suffers from issues such as incomplete topic coverage, misalignment of topics, and inefficiency. To address these limitations, we propose LLM-ITL, a novel LLM-in-the-loop framework that integrates LLMs with many existing Neural Topic Models (NTMs). In LLM-ITL, global topics and document representations are learned through the NTM, while an LLM refines the topics via a confidence-weighted Optimal Transport (OT)-based alignment objective. This process enhances the interpretability and coherence of the learned topics, while maintaining the efficiency of NTMs. Extensive experiments demonstrate that LLM-ITL can help NTMs significantly improve their topic interpretability while maintaining the quality of document representation.
@@ 1 Introduction
Topic modeling is an essential task in natural language processing that uncovers hidden thematic structures within large text collections in an unsupervised way. The ability to automatically extract topics has proven to be invaluable across a range of disciplines, such as bioinformatics (Liu et al., 2016), marketing research (Reisenbichler & Reutterer, 2019), and information retrieval (Yi & Allan, 2009). Topic models are usually based on probabilistic frameworks that generate a set of interpretable global topics, each represented as a distribution over vocabulary terms. These topics are then used to represent individual documents as mixtures of topics, providing a structured and interpretable view of the corpus. Recently, research on topic modeling has shifted from classical Bayesian methods such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to Neural Topic Models (NTMs) (Zhao et al., 2021; Churchill & Singh, 2022; Wu et al., 2024) that use deep neural networks to model document-topic distributions, enabling more expressive and flexible representations compared to their probabilistic counterparts.
@@
While Large Language Models (LLMs) (OpenAI, 2022; Touvron et al., 2023a; b) have redefined the landscape of natural language processing, topic models continue to hold their place as valuable tools for text analysis. Specifically, LLMs can provide a fine-grained understanding of a document; however, given a large collection of domain-specific documents, topic models are more suitable to obtain a clear global view of the topics in a more interpretable way with much less computational cost. Unsurprisingly, it has been a trending research direction to use LLMs to improve topic modeling (Rijcken et al., 2023; Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024; Chang et al., 2024). Despite the promising performance of these initial studies, most existing methods involve prompting LLMs to generate topics for each document in the corpus, which may lead to several limitations. As LLMs are asked to focus on a document individually, they may be unable to cover all the topics across all the documents in the corpus (Doi et al., 2024), which is critical in topic modeling. Moreover, although LLMs excel at capturing local context, they usually struggle with long documents with multiple interrelated topics, which may evolve or shift throughout the text. With their limited window of focus, LLMs may miss key topics of a document that are necessary to fully understand its content. Finally, it is computationally expensive as LLMs have to do inference for documents in the corpus; thus, existing methods usually scale poorly with large datasets.
@@
To overcome the aforementioned limitations, we propose LLM-ITL, a framework that integrates LLMs into NTMs and enhances the overall quality and interpretability of the learned topics, while maintaining computational efficiency. Specifically, to enhance the interpretability of the topics learned by an NTM, we introduce an LLM-based refinement step. The representative words for each topic, as generated by the NTM, are provided to the LLM, which suggests improved words that better capture the semantic meaning of the topic. The refinement process is guided by a novel plug-in objective based on Optimal Transport (OT), which ensures that the topics learned by the NTM align closely with the LLM’s refinements. Additionally, to mitigate potential hallucinations from the LLM (i.e., the generation of inaccurate or irrelevant suggestions), we introduce a confidence-weighted mechanism that adjusts the influence of the LLM’s suggestions based on their confidence scores. Our proposed LLM-ITL framework offers the following key contributions:
@@
  • Improved balance between topic coherence and document representation quality: With the LLM’s refined topics and OT-based alignment, the topics generated are more interpretable and semantically coherent. At the same time, LLM-ITL ensures that the document-topic distributions, as learned by the NTM, remain high-quality and reflective of the document’s content.
@@
  • Efficiency and scalability: Unlike most existing LLM-based approaches that rely on document-level LLM analysis, LLM-ITL uses LLMs at the word level, significantly reducing computational overhead for large datasets.
@@
  • State-of-the-art performance: Extensive experimental results on multiple benchmark datasets show that LLM-ITL achieves state-of-the-art performance in both topic coherence and document representation quality.
@@ 2.1 Problem Setup for Topic Modeling
Topic models have been popular across various fields for their ability to interpret text corpora in an unsupervised manner. Given a document collection $\mathcal{D}:=\{\bm{d}_{1},\ldots,\bm{d}_{N}\}$, a topic model learns to discover a set of global topics $\mathcal{T}:=\{\bm{t}_{1},\ldots,\bm{t}_{K}\}$, each of which is a distribution over the $V$ vocabulary words $\bm{t}\in\Delta^{V}$ ($\Delta$ denotes the probability simplex). Ideally, each topic represents a semantic concept that can be interpreted with its top-weighted words. At the document level, the topic model represents each document as a distribution over the $K$ topics, i.e., $\bm{z}\in\Delta^{K}$, which indicates the topic proportion of each topic within the document. The interpretability of topic models derives from both the corpus-level topics $\mathcal{T}$, and the document-level topical representation $\bm{z}$ for each document.
@@ 2.2 Neural Topic Models
A Neural Topic Model (NTM) (Miao et al., 2017; Srivastava & Sutton, 2017; Card et al., 2017; Dieng et al., 2020; Zhao et al., 2020; Xu et al., 2023a; b) is typically trained by modeling $p(\bm{z}|\bm{x})$ and $p(\bm{x}|\bm{z})$, where $\bm{x}\in\mathbb{N}^{V}$ represents the Bag-of-Words (BOWs) of a document. NTMs, which employ deep neural networks for topic modeling, are commonly based on Variational Auto-Encoders (VAEs) (Kingma & Welling, 2013) and Amortized Variational Inference (AVI) (Rezende et al., 2014). For VAE-NTMs, $p(\bm{x}|\bm{z})$ is modeled by a decoder network $\bm{\phi}$, i.e., $\bm{x}:=\bm{\phi}(\bm{z})$. The posterior $p(\bm{z}|\bm{x})$ is approximated by $q(\bm{z}|\bm{x})$, which is modeled by an encoder network $\bm{\theta}$, i.e., $\bm{z}:=\bm{\theta}(\bm{x})$. The training objective of VAE-NTMs is to maximize the Evidence Lower Bound (ELBO):
$\max_{\bm{\theta},\bm{\phi}}{(\mathbb{E}_{q_{\bm{\theta}}(\bm{z}|\bm{x})}[\log{p_{\bm{\phi}}(\bm{x}|\bm{z})}]}-\mathbb{KL}[q_{\bm{\theta}}(\bm{z}|\bm{x})\parallel p(\bm{z})]),$ (1)
@@ 2.3 Optimal Transport
Figure 1: Overview of LLM-ITL. The topics and document representations are learned by the Neural Topic Model (NTM) component. Concurrently, a Large Language Model (LLM) suggests better topic words for the learned topics (e.g., decoder) from the NTM. An Optimal Transport (OT)-based topic alignment objective is proposed to align the word distribution between the topics from the NTM and those suggested by the LLM. The alignment is further weighted by the confidence of the LLM in providing the suggestions. This confidence-weighted topic refinement objective is plugged into the standard training of an NTM as the overall objective of LLM-ITL.
@@ 3.1 LLM-based Topic Suggestion
where $\bm{w}$ represents the original topic words; $\bm{\theta}^{\text{llm}}$ denotes the LLM model; the topic label $\bm{w}^{l}$ and the refined words $\bm{w}^{\prime}$ are extracted as subsequences from the LLM’s output $\bm{s}$. The used prompt is illustrated in Appendix A.1. A study of prompt variants is provided in Appendix H.
@@ 3.2 OT-based Topic Alignment
A key innovation in LLM-ITL is the use of Optimal Transport (OT) distance to align the topic word distributions generated by the NTM with the refined topic word distributions provided by the LLM. OT is a mathematical framework that computes the “cost” of transforming one probability distribution into another, making it an ideal tool for measuring the alignment between two sets of words (Kusner et al., 2015; Yang et al., 2024).
@@ 3.3 Confidence-Weighted Topic Refinement
We propose two methods for calculating topic labeling confidence, considering whether the LLM is open-source or not: (1) Label token probability, applicable for open-source LLMs where the token probability of their generation is accessible; (2) Word intrusion confidence, available for both open and closed-source LLMs.
@@
This method evaluates the proportion of irrelevant or “intruder” words removed by the LLM during suggestion. A topic label generated based on a higher rate of intruder removal indicates that it is harder for the LLM to identify the topic from the original topic words, leading to lower confidence:
@@
By incorporating the topic labeling confidence as a weight for the topic alignment loss, we adaptively adjust the impact of the LLM’s suggestion based on the confidence score. We write our confidence-weighted topic refinement objective as follows:
@@
3.4 Integration with NTMs
One of the core strengths of LLM-ITL lies in its flexibility to integrate with various NTMs while leveraging the semantic capabilities of LLMs for topic refinement. The framework is designed to complement and enhance NTMs, providing an efficient and interpretable topic modeling pipeline. The LLM-ITL framework is highly modular and can be seamlessly integrated with a wide range of NTMs. Here, we focus on the VAE-NTM framework that accommodate many NTMs (Miao et al., 2017; Srivastava & Sutton, 2017; Card et al., 2017; Dieng et al., 2020; Zhao et al., 2020; Nguyen & Luu, 2021; Xu et al., 2023a; b), while our framework is not limited to VAE-NTMs only. By integrating the topic refinement objective with the training of an NTM, we obtain the overall objective of LLM-ITL:
@@ 3.4 Integration with NTMs
where $\bm{\Theta}:=\{\bm{\theta},\bm{\phi}\}$ denotes model parameters; $\mathcal{L}^{\text{ntm}}$ and $\mathcal{L}^{\text{refine}}$ denote the NTM loss and refinement loss in Eq. 1 and Eq. 10, respectively; $\gamma$ controls the strength of focusing on the LLM’s refinements; $t$ and $T^{\text{refine}}$ denote the current training step and the start of topic refinement, respectively; $\mathbf{I}(\cdot)$ denotes the indicator function, which ensures that the refinement process only starts after the NTM has learned a stable topic representation, allowing the model to capture the core structure of the corpus before fine-tuning the topics with LLM guidance. The algorithm of LLM-ITL is provided in Appendix B.
@@ 4 Related Work
Classical topic models, such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and its variants (Blei & Lafferty, 2006; Rosen-Zvi et al., 2012; Yan et al., 2013), are Bayesian probabilistic models with various generative assumptions about the documents. Neural Topic Models (NTMs) (Miao et al., 2017; Srivastava & Sutton, 2017; Card et al., 2017; Dieng et al., 2020; Zhao et al., 2020; Nguyen & Luu, 2021; Xu et al., 2023a; b) use deep neural networks to learn topics and document representations, and are commonly based on Variational Autoencoders (VAE) (Kingma & Welling, 2013) and Amortized Variational Inference (AVI) (Rezende et al., 2014). Clustering-based topic models (Sia et al., 2020; Grootendorst, 2022) discover topics using clustering algorithms based on embeddings from pre-trained language models. Ultimately, the capability of these models to interpret a corpus is limited by the top words representation of each topic.
@@
LLMs have been involved in topic modeling in various ways. Rijcken et al. (2023) investigate the use of ChatGPT (OpenAI, 2022) to generate descriptions for topic words and found the effectiveness of these topic descriptions. Recent works leverage LLMs for topic model evaluation in different ways, such as applying LLMs for word intrusion or topic rating for topics (Rahimi et al., 2023; Stammbach et al., 2023), or keyword generation for documents (Yang et al., 2024). LLM-based topic models have emerged (Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024), which prompt LLMs to generate topics and assign topics to documents. Different from these methods that focus on the document-level, ours prompts LLMs to suggest better topic words which are used to refine the training of NTMs. More recently, Chang et al. (2024) show that LLMs are effective at refining topic words, leading to improved topic coherence. However, their method refines the topic words of trained topic models in a post-hoc manner, while ours is a regularization term for training NTMs. Our method is also loosely related to uncertainty estimation of LLMs and we omit the discussion on this in Appendix C.

### v2

@@ Abstract
Topic modeling is a fundamental task in natural language processing, allowing the discovery of latent thematic structures in text corpora. While Large Language Models (LLMs) have demonstrated promising capabilities in topic discovery, their direct application to topic modeling suffers from issues such as incomplete topic coverage, misalignment of topics, and inefficiency. To address these limitations, we propose LLM-ITL, a novel LLM-in-the-loop framework that integrates LLMs with Neural Topic Models (NTMs). In LLM-ITL, global topics and document representations are learned through the NTM. Meanwhile, an LLM refines these topics using an Optimal Transport (OT)-based alignment objective, where the refinement is dynamically adjusted based on the LLM’s confidence in suggesting topical words for each set of input words. With the flexibility of being integrated into many existing NTMs, the proposed approach enhances the interpretability of topics while preserving the efficiency of NTMs in learning topics and document representations. Extensive experiments demonstrate that LLM-ITL helps NTMs significantly improve their topic interpretability while maintaining the quality of document representation. Our code and datasets will be available at Github.
Neural Topic Modeling with Large Language Models in the Loop
Xiaohao Yang1, He Zhao2, Weijie Xu3, Yuanyuan Qi1, Jueqing Lu1, Dinh Phung1, Lan Du1 1Faculty of IT, Monash University, Australia 2CSIRO’s Data61, Australia  3Amazon, America {xiaohao.yang,yuanyuan.qi,jueqing.lu,dinh.phung,lan.du}@monash.edu he.zhao@data61.csiro.au,  weijie.xu@amazon.com
@@ 1 Introduction
Topic modeling is an essential task in natural language processing that uncovers hidden thematic structures within text collections in an unsupervised way. The ability to automatically extract topics has proven to be invaluable across a range of disciplines, such as bioinformatics (Liu et al., 2016), marketing research (Reisenbichler and Reutterer, 2019), and information retrieval (Yi and Allan, 2009). Topic models are conventionally based on probabilistic frameworks that generate a set of interpretable global topics, each represented as a distribution over vocabulary terms as Latent Dirichlet Allocation (LDA) (Blei et al., 2003). These topics are then used to represent individual documents as mixtures of topics, providing a structured and interpretable view of the corpus. Recently, research on Neural Topic Models (NTMs) (Zhao et al., 2021; Churchill and Singh, 2022; Abdelrazek et al., 2023; Wu et al., 2024a) has been popular, which uses deep neural networks to model document-topic distributions, enabling more expressive and flexible representations compared to their probabilistic counterparts.
@@
While Large Language Models (LLMs) (OpenAI, 2022; Touvron et al., 2023a, b) have redefined the landscape of natural language processing, topic models can still be valuable tools for text analysis. Specifically, LLMs can provide a fine-grained understanding of a document; however, given a large collection of domain-specific documents, topic models are more suitable to obtain a clear global view of the topics in a more interpretable way with much less computational cost. Unsurprisingly, it has been a trending research direction to use LLMs to improve topic modeling (Rijcken et al., 2023; Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024; Chang et al., 2024). Despite the promising performance of these initial studies, most existing methods involve prompting LLMs to generate topics for each document in the corpus, which may lead to several limitations. As LLMs are asked to focus on a document individually, they may be unable to cover global topics across all the documents in the corpus (Doi et al., 2024), which is critical in topic modeling. Moreover, although LLMs excel at capturing local context, they usually struggle with long documents with multiple interrelated topics, which may evolve or shift throughout the text. With their limited window of focus, LLMs may miss key topics of a document that are necessary to fully understand its content. Finally, it is computationally expensive as LLMs have to do inference for every document in the corpus or subsets; thus, existing methods usually scale poorly with large datasets.
@@
To overcome the aforementioned limitations, we propose LLM-ITL, a framework that integrates LLMs into NTMs and enhances the overall quality and interpretability of the learned topics, while maintaining computational efficiency. Specifically, to enhance the interpretability of the topics learned by an NTM, we introduce an LLM-based refinement step. The representative words for each topic, as generated by the NTM, are provided to the LLM, which suggests improved words that better describe the semantic meaning of the topic. The refinement process is guided by a novel plug-in objective based on Optimal Transport (OT), which ensures that the topics learned by the NTM align closely with the LLM’s refinements. Additionally, to mitigate potential hallucinations from the LLM (i.e., the generation of inaccurate or irrelevant suggestions) for hard or incoherent topics, we introduce a confidence-weighted mechanism that adjusts the influence of the LLM’s suggestions based on their confidence scores. Finally, to balance learning from the input corpus by the NTM and refinement by the LLM, a warm-up phase is adopted to ensure that the NTM component learns corpus-relevant topics and document representations before the LLM’s refinement is applied. Our proposed LLM-ITL framework offers the following key contributions:
@@
  • Boosting topic quality while maintaining document representation quality: With the LLM’s refined topics and OT-based alignment, the topics generated are more interpretable and semantically coherent. At the same time, LLM-ITL ensures that the document-topic distributions, as learned by the NTM, remain high-quality and reflective of the document’s content.
@@
  • Efficiency in using LLMs: Unlike most existing LLM-based approaches that rely on document-level LLM analysis, LLM-ITL uses LLMs at the word level, significantly reducing computational overhead.
@@
Figure 1: LLM-ITL overview. The topics and document representations are learned by the Neural Topic Model (NTM) component. After a warm-up stage, a Large Language Model (LLM) suggests better topic words for the learned topics (e.g., decoder) from the NTM. An Optimal Transport (OT)-based topic alignment objective is proposed to align the word distribution between the topics from the NTM and those suggested by the LLM. The alignment is further weighted by the confidence of the LLM in providing the suggestions. This confidence-weighted topic refinement objective is plugged into the standard training of an NTM as the overall objective of LLM-ITL.
@@ 2.1 Problem Setup for Topic Modeling
Given a document collection $\mathcal{D}:=\{\bm{d}_{1},\ldots,\bm{d}_{N}\}$, a topic model learns to discover a set of global topics $\mathcal{T}:=\{\bm{t}_{1},\ldots,\bm{t}_{K}\}$, each of which is a distribution over the $V$ vocabulary words $\bm{t}\in\Delta^{V}$ ($\Delta$ denotes the probability simplex). Ideally, each topic represents a semantic concept that can be interpreted with its top-weighted words. At the document level, the topic model represents each document as a distribution over the $K$ topics, i.e., $\bm{z}\in\Delta^{K}$, which indicates the topic proportion of each topic within the document. The interpretability of topic models derives from both the corpus-level topics $\mathcal{T}$, and the document-level topical representation $\bm{z}$ for each document.
@@ 2.2 Neural Topic Models
A Neural Topic Model (NTM) is typically trained by modeling $p(\bm{z}|\bm{x})$ and $p(\bm{x}|\bm{z})$, where $\bm{x}\in\mathbb{N}^{V}$ represents the Bag-of-Words (BOWs) of a document. NTMs, which employ deep neural networks for topic modeling, are commonly based on Variational Auto-Encoders (VAEs) (Kingma and Welling, 2013) and Amortized Variational Inference (AVI) (Rezende et al., 2014). For VAE-NTMs, $p(\bm{x}|\bm{z})$ is modeled by a decoder network $\bm{\phi}$, i.e., $\bm{x}:=\bm{\phi}(\bm{z})$. The posterior $p(\bm{z}|\bm{x})$ is approximated by $q(\bm{z}|\bm{x})$, which is modeled by an encoder network $\bm{\theta}$, i.e., $\bm{z}:=\bm{\theta}(\bm{x})$. The training objective of VAE-NTMs is to maximize the Evidence Lower Bound (ELBO):
$\max\limits_{\bm{\theta},\bm{\phi}}{(\mathbb{E}_{q_{\bm{\theta}}(\bm{z}|\bm{x})}[\log{p_{\bm{\phi}}(\bm{x}|\bm{z})}]}-\mathbb{KL}[q_{\bm{\theta}}(\bm{z}|\bm{x})\parallel p(\bm{z})]),$
@@ 3.1 LLM-based Topic Suggestion
where $\bm{w}$ represents the original topic words; $\bm{\theta}^{\text{llm}}$ denotes the LLM model; the topic label $\bm{w}^{l}$ and the refined words $\bm{w}^{\prime}$ are extracted as subsequences from the LLM’s output $\bm{s}$. Notably, the words in $\bm{w}^{\prime}$ are further filtered based on the corpus vocabulary to ensure that no out-of-vocabulary (OOV) words are included. Our used prompt is illustrated in Figure A1. A study of prompt variants is provided in Appendix H.
@@ 3.2 OT-based Topic Alignment
A key innovation in LLM-ITL is the use of Optimal Transport (OT) distance to align the topic word distributions generated by the NTM with the refined topic word distributions provided by the LLM. OT is a mathematical framework that computes the “cost” of transforming one probability distribution into another, and has shown its effectiveness in measuring the alignment between two sets of words (Kusner et al., 2015; Yang et al., 2024).
@@ 3.3 Confidence-Weighted Topic Refinement
We propose two methods for calculating topic labeling confidence of the LLM, considering whether the LLM is open-source or not: (1) Label token probability, applicable for open-source LLMs where the token probability of their generation is accessible; (2) Word intrusion confidence, available for both open and closed-source LLMs.
@@
This method evaluates the proportion of irrelevant or “intruder" words removed by the LLM during suggestion. A topic label generated based on a higher rate of intruder removal indicates that it is harder for the LLM to identify the topic from the original topic words, leading to lower confidence:
@@
By incorporating the topic labeling confidence (Eq. 8 or Eq. 9) as a weight for the topic alignment loss, we adaptively adjust the impact of the LLM’s suggestion based on the confidence score. We write our confidence-weighted topic refinement objective as follows:
@@
3.4 Integration with Warm-Up
While the LLM’s refinement can lead to more coherent topics, over-reliance on it may bias the topics toward the LLM’s knowledge rather than reflecting the global information of the input corpus, thereby harming the topical representation of documents. To address this issue, we propose integrating the NTM with the LLM after a warm-up stage. This ensures that the refinement process begins only after the NTM has learned a stable topic representation, allowing the model to capture the core structure of the corpus before fine-tuning the topics with LLM guidance. By integrating the topic refinement objective with the training of an NTM with warm-up, we obtain the overall objective of LLM-ITL:
@@ 3.4 Integration with Warm-Up
where $\bm{\Theta}:=\{\bm{\theta},\bm{\phi}\}$ denotes model parameters; $\mathcal{L}^{\text{ntm}}$ and $\mathcal{L}^{\text{refine}}$ denote the NTM loss and refinement loss in Eq. 1 and Eq. 10, respectively; $\gamma$ controls the strength of focusing on the LLM’s refinements; $t$ and $T^{\text{refine}}$ denote the current training step and the start of topic refinement, respectively; $\mathbf{I}(\cdot)$ denotes the indicator function. We provide the algorithm of LLM-ITL in Appendix B.
@@ 4 Related Work
Classical topic models, such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and its variants (Blei and Lafferty, 2006; Rosen-Zvi et al., 2012; Yan et al., 2013), are Bayesian probabilistic models with various generative assumptions about the documents. Neural Topic Models (NTMs) (Miao et al., 2017; Srivastava and Sutton, 2017; Card et al., 2017; Dieng et al., 2020; Bianchi et al., 2020; Zhao et al., 2020; Nguyen and Luu, 2021; Shen et al., 2021; Duan et al., 2021; Li et al., 2022; Xu et al., 2023a, b; Yang et al., 2023; Miyamoto et al., 2023; Wu et al., 2024b; Chen et al., 2025a, b) use deep neural networks to learn topics and document representations. Clustering-based topic models (Sia et al., 2020; Grootendorst, 2022; Angelov and Inkpen, 2024) discover topics using clustering algorithms based on embeddings from pre-trained language models. Ours is not a specific topic model, but rather a general framework designed to integrate LLMs with a wide range of NTMs.
@@
LLMs have been involved in topic modeling in various ways. Rijcken et al. (2023) investigate the use of ChatGPT (OpenAI, 2022) to generate descriptions for topic words and found the effectiveness of these topic descriptions. Recent works leverage LLMs for topic model evaluation in different ways, such as applying LLMs for word intrusion or topic rating for topics (Rahimi et al., 2023; Stammbach et al., 2023), or keyword generation for documents (Yang et al., 2024). LLM-based topic models have emerged (Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024), which prompt LLMs to generate topics and assign topics to documents. Unlike existing methods that focus on document-level analysis relying on LLMs, our approach prompts LLMs to analyze a group of topic words, which are then used to guide the training of NTMs. More recently, Chang et al. (2024) show that LLMs are effective at refining topic words, leading to improved topic coherence. However, their method refines the topic words of trained topic models in a post-hoc manner, whereas ours incorporates refinement as a regularization term during training. Our method is also loosely related to uncertainty estimation of LLMs, and we omit the discussion on this in Appendix C.